# Driving with Prior Maps: Unified Vector Prior Encoding for Autonomous Vehicle Mapping

### v2

@@ Introduction
We mainly validate our method on the popular nuScenes (Caesar et al. 2019) dataset following the standard setting of previous methods (Li et al. 2021; Liu et al. 2022; Liao et al. 2022). NuScenes is a large-scale autonomous driving benchmark, that includes multiple traversals with precise localization and annotated HD maps labels. It contains a total of 1,000 scenes, each with an approximate duration of 20 seconds. The data was collected utilizing a 32-beam LiDAR operating at 20Hz and six cameras offering a 360-degree field of view at 12 Hz. Annotations for keyframes are furnished at a rate of 2Hz, with the training and validation sets containing 28,130 frames and 6,019 frames, respectively.
@@
For the vectorized results, we follow the standard metric used in previous works (Li et al. 2021; Liu et al. 2022; Liao et al. 2022). The perception range along the direction of vehicle travel are $[-15.0m,15.0m]$ for the $X$-axis and $[-30.0m,30.0m]$ for the $Y$-axis. And we adopt average precision (AP) evaluated the quality of map construction. Chamfer distance $D_{Chamfer}$ is utilized to determine whether the prediction and ground truth (GT) are matched or not. We compute the $\mathrm{AP}_{\tau}$ under several $D_{Chamfer}$ thresholds $(\tau\in T,T=\{0.5,1.0,1.5\})$, and then calculate the mean across all thresholds as the the final AP metric:
@@
Following the previous methods (Li et al. 2021; Liu et al. 2022; Liao et al. 2022), we have selected three kinds of map elements are chosen for fair evaluation: lane divider, pedestrian crossing and road boundary. Furthermore, to evaluate the effectiveness of our method across different evaluation metrics and for the rasterized results, we follow the previous works (Li et al. 2021; Xiong et al. 2023; Jiang et al. 2024) and employ intersection-over-union (IoU) as the metrics for segmentation results.
@@
We primarily perform our experiments using the latest and influential vectorized model MapTRv2 (Liao et al. 2023). In order to validate the broad applicability of our method, we also use HDMapNet (Li et al. 2021) as baseline. For fair comparisons, experiments were conducted on all baseline models using their original configs and parameters without any changes. All experiments were trained with 6 NVIDIA RTX A6000. For pre-training UVE, we trained the 24-epoch on nuScenes vecror data. For historical prediction maps, we run each base model to get their respective online local maps.
HDMapNet C Effi-B0 30 IoU 18.7 40.6 39.5 32.9 -
+ SD map Prior C Effi-B0 30 IoU 20.8 41.7 41.8 34.8 -
+ Online Local Map Prior C Effi-B0 30 IoU 22.5 41.3 41.7 35.1 -
+ HD map-EX* Prior C Effi-B0 30 IoU 21.2 42.3 44.9 36.1 -
MapTRv2 C R50 24 AP 59.8 62.4 62.4 61.5 38.6
+ SD map Prior C R50 24 AP 61.7 65.5 66.4 64.5 42.7
+ Online Local Map Prior C R50 24 AP 63.6 67.3 66.2 65.7 44.7
+ HD map-EX* Prior C R50 24 AP 61.5 65.6 68.3 65.1 44.3
Table 1: Performance of various online mapping models with different prior maps on nuScenes within a $60m\times 30m$ perception range. “C” denotes camera input. The best results using the same backbone are highlighted in bold. The SD map is sourced from OpenStreetMap (Haklay and Weber 2008). The online local map refers to the historical prediction results. As the nuScenes dataset lacks existing HD maps, we followed the approach of MapEX (Sun et al. 2023) to create an HD map-EX*, simulating existing HD maps by removing pedestrian crossings and lane dividers, while retaining only the road boundaries.
@@
Table 3: Ablation experiments of the pre-training strategies on MapTRv2. Here, we fixed prior_num=2, search_range=5m and integration=replace. “Seg” and “Pt” represent segment-level and point-level, respectively.
Main Results
To evaluate the effectiveness of our method across different model architectures, evaluation metrics, and vector prior maps, we integrated the UVE-encoded prior maps into both HDMapNet and the vectorized model MapTRv2 (see Table 1). The prior information provided by SD maps, which primarily consists of a centerline skeleton, resulted in the smallest improvement. Specifically, MapTRv2 and HDMapNet showed enhancements of 3.0 mAP and 1.9 mIoU, respectively, when using SD map priors. And for query-based model, our method has a greater improvement because our HPQuery makes better use of prior information. The incorporation of online local maps, which offer up-to-date information, led to more significant improvements. This is reflected in the enhancements of 4.2 mAP and 2.2 mIoU for MapTRv2 and HDMapNet, respectively. These results indicate that more accurate and current local maps contribute to better performance by providing more comprehensive and precise information. Figure 4 illustrates how online local maps effectively restore obscured lane dividers, pedestrian crossings, and road boundaries. Although offline HD maps are highly accurate, their lack of complete prior information resulted in notable improvements of 3.6 mAP and 3.2 mIoU for MapTRv2 and HDMapNet, respectively. Additionally, we evaluated the results using a stricter threshold of 0.5m. Under this criterion, our method with SD map, online local map, and outdated HD map (HD map-EX*) priors achieved improvements of 4.1 mAP, 6.1 mAP, and 5.7 mAP, respectively. These results underscore that our method provides more substantial improvements under stricter evaluation conditions. Overall, our method demonstrated consistent performance improvements across all baseline models with various prior maps, highlighting its generalizability and potential applicability to other mapping frameworks.
Ablation Study
@@
The Effect of Pre-training.
To effectively capture the prior information of vector maps, we pre-trained UVE using different strategies, as shown in Table 3. When we fed the original vector map directly to UVE without any noise, UVE converged with a minimal error of 0.09 meters. However, the simplicity of the task might have limited the model’s ability to fully learn the prior distribution, resulting in only a 0.3 mAP improvement. We added segment-level Gaussian noise with a 3-meter standard deviation to 20% of the map elements randomly. The UVE’s average Euclidean distance error in reconstructing the entire map was 0.46 meters. However, this approach led to a slight decrease in mAP, likely because the UVE had not fully converged. Next, we applied 1-meter segment-level noise to 10% of the map elements randomly, resulting in a reduced error of 0.15 meters for UVE and a 0.2 mAP improvement compared to the model without pre-training. Building on this, we introduced 1-meter point-level noise to 5% of the vector points. Although this increased UVE’s error to 0.18 meters, it yielded a more substantial mAP improvement of 0.6. This demonstrates that pre-training strategies incorporating both segment-level and point-level noise are more effective at learning the prior information of vector maps. We also experimented with masking instead of noise, which increased UVE’s error to 0.43 meters but still resulted in a 0.4 mAP improvement. This suggests that while challenging tasks based on masking can pre-train UVE effectively, their impact may be constrained by the amount of pre-training data.
Impact of the Number of Prior Maps.
We conducted an ablation study to assess the effect of varying the number of prior maps on model performance (see Table 4). The results indicate that using a single prior map led to a 0.9 mAP improvement, demonstrating its effectiveness in enhancing current perception. Incorporating two prior maps resulted in a more substantial improvement of 1.5 mAP, reflecting the value of richer prior information. However, using multiple prior maps only yielded a 1.3 mAP improvement. This could be attributed to the integration method, where replacing too many original queries with prior features may have diminished the performance of the original queries.
Effect of Different Search Ranges for Prior Maps.
Different Methods for Integrating Prior Maps.
To evaluate the versatility of our approach, we tested various integration methods in Table 6. Compared to MapTRv2, the HPQuery methods of replacement and addition resulted in mAP improvements of 2.1 and 1.6, respectively, but these methods may impact the performance of the original query. In contrast, the concatenation method achieved a more substantial improvement of 4.2 mAP. This approach effectively preserves the original query’s integrity while also leveraging the instance-level and point-level prior features learned by UVE, making it the most effective integration method.
Inference Speed, Memory, Model Size, and Detailed Runtime.
@@
search_range=5m One 61.9 62.1 62.4 62.4
pretrain=false Two 61.3 63.4 64.4 63.0
integration=replace Multiple 61.2 63.0 64.2 62.8
@@
prior_num=2 5 m 61.3 63.4 64.4 63.0
pretrain=false 10 m 60.8 63.7 64.0 62.8
integration=replace 15 m 61.5 62.7 64.0 62.7
@@
prior_num=2 Replace 61.6 64.3 64.8 63.6
pretrain=true Add 62.5 63.3 63.6 63.1
@@
Table 6: Ablation experiments on the integrated methods of online local prior maps on MapTRv2. Note that we proposed other integrated methods after pre-training UVE, so we use pre-trained UVE here.
@@
Figure 4: Qualitative results with and w/o online local prior. Prior maps can help restore obscured map elements and make predictions more complete and accurate.
Conclusion
In this paper, we introduced PriorDrive, a novel framework that effectively leverages various types of prior maps to enhance the accuracy and robustness of online HD map construction for autonomous vehicles. Central to our approach is the UVE, which we designed to encode diverse vector data efficiently. Through comprehensive experimentation, we demonstrated that UVE, combined with our proposed pre-training strategies, significantly improves the performance of state-of-the-art mapping models. Our approach not only addresses the challenges associated with real-time HD map construction in dynamic and complex environments but also offers a scalable solution that continually improves map accuracy over time. The iterative use of historical prediction maps as priors leads to progressively refined map outputs, thus ensuring that the most up-to-date road information is available for autonomous navigation.

### v3

@@ 4.1 Experimental settings
We mainly validate our method on the popular nuScenes [1] and Argoverse 2 [33] following the standard setting of previous methods [25, 19]. NuScenes contains a total of 1,000 scenes, each with an approximate duration of 20 seconds. The data was collected utilizing a 32-beam LiDAR operating and six cameras offering a 360-degree field of view. Argoverse 2 includes 1000 logs, each offering 15 seconds of 20Hz RGB images captured from 7 cameras, along with a 3D vector map.
Furthermore, to demonstrate that PriorDrive is plug-and-play and widely applicable to other vector tasks, we conducted a topology inference task on OpenLane-V2 [31] dataset. Subset A, developed by Argogorse 2 [33], includes seven views as input, along with an additional SD map.
@@
For the vectorized results, we follow the standard metric used in previous works [15, 25, 19]. The perception range along the direction of vehicle travel are $[-15.0m,15.0m]$ for the $X$-axis and $[-30.0m,30.0m]$ for the $Y$-axis. And we adopt average precision (AP) evaluated the quality of map construction. We compute the $\mathrm{AP}_{\tau}$ under several Chamfer distance $D_{Chamfer}$ thresholds $(\tau\in T,T=\{0.5,1.0,1.5\})$, and then calculate the mean across all thresholds as the the final AP metric:
@@
Furthermore, to evaluate the effectiveness of our method across different evaluation metrics and for the rasterized results, we follow the previous works [15, 35, 13] and employ IoU as the metrics for segmentation results.
For OpenLane-V2’s lane centerline perception task, $DET_{l}$ averages the Frechet distance to quantify similarity. $DET_{t}$ uses IoU and averages it across various traffic categories. $TOP_{ll}$ and $TOP_{lt}$ compute the topology matrix similarity between lanes and between lanes and traffic elements, respectively. The overall metric is called OLS:
$\mathrm{OLS}=\frac{1}{4}[\mathrm{DET}_{l}+\mathrm{DET}_{t}+f(\mathrm{TOP}_{ll})+f(\mathrm{TOP}_{lt})],$ (12)
where $f$ is the square root function.
@@
For fair comparisons, experiments were conducted on all baseline models using their original configs and parameters without any changes. All experiments were trained with 6 NVIDIA RTX A6000. For pre-training UVE, we trained the 24-epoch on nuScenes vecror data. For historical prediction maps, we run each base model to get their respective online local maps.
Method $\textbf{AP}_{ped}$ $\textbf{AP}_{div}$ $\textbf{AP}_{bou}$ mAP
MGMap [22] 52.8 67.5 68.1 62.8
MapQR [26] 60.1 71.2 66.2 65.9
MapTRv2 [20] 60.7 68.9 64.5 64.7
# SD Prior (Ours) 62.0 70.2 66.3 66.2
# Online Local Prior (Ours) 63.9 71.7 68.6 68.1
# HD map-EX* (Ours) 61.5 69.8 69.4 66.9
Table 2: Comparison on Argoverse 2 where the 3D map elements are predicted directly.
4.2 Main results
Results on nuScenes.
To evaluate the effectiveness of our method across different model, evaluation metrics, and vector prior maps, we choose HDMapNet and MapTRv2 as baselines (see Table 1). SD maps, which primarily consists of a centerline skeleton, showed enhancements of 3.0 mAP for MapTRv2 and 2.3 mIoU for HDMapNet, respectively. And for query-based model, our method has a greater improvement because our HPQuery makes better use of prior maps. Online local maps, which offer up-to-date information, led to more significant improvements, with 4.2 mAP and 2.6 mIoU for MapTRv2 and HDMapNet, respectively. These results indicate that more accurate and current local maps contribute to better performance by providing more comprehensive and precise information. Figure 5 illustrates how online local maps effectively restore obscured map elements. More visualization are shown in the supplementary material. Although offline HD maps are highly accurate, their lack of complete prior information resulted in notable improvements of 3.6 mAP and 3.2 mIoU for MapTRv2 and HDMapNet, respectively. Additionally, we evaluated the results using a stricter threshold of 0.5m, achieving improvements of 4.1 mAP, 6.1 mAP, and 5.7 mAP, respectively, underscoring that our method provides more substantial improvements under stricter evaluation conditions. Overall, our method demonstrated consistent performance improvements across all baselines with various prior maps, highlighting its generalizability and potential applicability to other mapping frameworks.
Method $\textbf{DET}_{l}$ $\textbf{DET}_{t}$ $\textbf{TOP}_{ll}$ $\textbf{TOP}_{lt}$ OLS
MapTR [19] 17.7 43.5 5.9 15.1 31.0
TopoNet [17] 28.6 48.6 10.9 23.8 39.8
TopoLogic [7] 29.9 47.2 23.9 25.4 44.1
# SD Prior (Ours) 29.9 47.1 29.6 32.9 47.2
# Online Local Prior (Ours) 31.0 47.8 30.9 32.5 47.9
Table 3: Performance comparison with SOTA methods on OpenLaneV2 subset_A set, where SD Prior comes from the SD map of the openLanev2 itself. Since the map elements here have only centerline and no boundaries, there is no HD map-EX* prior.
Results on Argoverse 2.
In Table 2, we present the results of Argoverse 2, which provides 3D vectorized map elements with additional height information as ground truth. These testings are all trained 6 epochs with ResNet50 as the backbone. Our PriorDrive improves performance by 1.5, 3.4, and 2.2, respectively, using different prior maps, demonstrating the superior generalization ability of PriorDrive in terms of 3D vectorized map construction.
Results on OpenLane-V2.
We compared PriorDrive with SOTA methods when using ResNet-50 for training 24 epochs. Table 3 shows the results on subset A of OpenLane-V2. Compared to TopoLogic, our method achieves better topological reasoning accuracy using connected SD priors (29.6 v.s. 23.9 on $TOP_{ll}$, 32.9 v.s. 25.4 on $TOP_{lt}$), and the overall score OLS increased by 3.1, indicating that connected SD is more beneficial for lane connectivity and correspondence with traffic elements. Employing online local priors resulted in decent detection accuracy for both detection and topological reasoning, with OLS rising to 47.9. Overall, PriorDrive is a plug-and-play solution that can be widely applied to other vector tasks.
@@ 4.2 Main results
Table 5: Ablation of the pre-training strategies. Here, we fixed prior_num=multiple, search_range=5m and integration=replace. “Seg” and “Pt” represent segment-level and point-level, respectively. * indicates data augmentation to expand the data volume.
4.3 Ablation study
@@
search_range=5m One 61.1 63.8 63.6 62.8
pretrain=true Two 61.4 66.2 63.0 63.5
integration=replace Multiple 62.5 66.3 64.5 64.4
@@
prior_num=2 5 m 62.5 66.3 64.5 64.4
pretrain=true 10 m 61.4 64.6 64.4 63.5
integration=replace 15 m 61.9 63.2 63.9 63.0
@@
The effect of pre-training.
To effectively capture the prior information of vector maps, we pre-trained UVE using different strategies in Table 5. When we fed the original vector map directly to UVE without any noise, UVE converged with a minimal error of 0.09 meters. However, the simple task might have limited the model’s ability to fully learn the prior distribution, resulting in only a 0.3 mAP improvement. We added segment-level Gaussian noise with a 3-meter standard deviation to 20% of the map elements randomly. The UVE’s average Euclidean distance error was 0.46 meters. However, this approach led to a slight decrease in mAP, likely because the UVE had not fully converged. Next, we applied 1-meter segment-level noise to 10% of the map elements randomly, resulting in a reduced error of 0.15 meters for UVE and a 0.2 mAP improvement compared to the model without pre-training. Building on this, we introduced 1-meter point-level noise to 5% of the vector points. Although this increased UVE’s error to 0.18 meters, it yielded a more substantial mAP improvement of 0.6. This demonstrates that pre-training strategies incorporating both segment-level and point-level noise are more effective at learning the prior information of vector maps. We also experimented with masking instead of noise, which increased UVE’s error to 0.43 meters but still resulted in a 0.4 mAP improvement. Although various proposed tasks can effectively pre-train UVE, they are still limited by the amount of pre-training data. Therefore, we employ horizontal and vertical flipping data augmentation to expand the pre-training dataset. The result of the data augmentation is shown as *, effectively improving 1.4 mAP.
@@
prior_num=2 Replace 62.5 66.3 64.5 64.4
pretrain=true Add 63.3 64.1 64.4 63.9
@@
Table 8: Ablation experiments on the integrated methods of online local prior maps on MapTRv2.
Impact of the number of prior maps.
We conducted an ablation study to assess the effect of the number of prior maps on model performance (see Table 6). The results indicate that using a single prior map led to a 1.3 mAP improvement, demonstrating its effectiveness in enhancing current perception. Incorporating two prior maps resulted in a more substantial improvement of 2.0 mAP, reflecting the value of richer prior information. Finally, multiple prior maps effectively improved 2.9 mAP, indicating that more prior maps can provide more comprehensive necessary information.
Effect of different search ranges for prior maps.
Different methods for integrating prior maps.
To evaluate the versatility of our approach, we tested various integration methods in Table 8. Compared to MapTRv2, the HPQuery methods of replacement and addition resulted in mAP improvements of 2.9 and 2.4, respectively, but these methods may impact the performance of the original query. In contrast, the concatenation method achieved a more substantial improvement of 4.2 mAP. This approach effectively preserves the original query’s integrity while also leveraging the instance-level and point-level prior features learned by UVE, making it the most effective integration method.
Inference speed, memory, model size, and detailed runtime.
@@
Figure 5: Qualitative results with and w/o online local prior.
5 Conclusion
In this paper, we introduced PriorDrive, a novel framework that effectively leverages various types of prior maps to enhance the accuracy and robustness of online HD map construction. Central to PriorDrive is UVE, which can encode diverse vector data efficiently. Through comprehensive experiments, we demonstrated that UVE, combined with proposed pre-training strategies, significantly improves performance of state-of-the-art models in various vector tasks. PriorDrive not only addresses the challenges associated with online HD map construction in complex environments but also offers a scalable solution that continually improves accuracy over time. The iterative use of historical prediction maps as priors leads to progressively refined map outputs, providing latest road information for autonomous driving.